OpenStudy (anonymous):

Calculate the power of i. i^-22

OpenStudy (anonymous):

i^0 = 1 i^1 = i i^2 = -1 i^3 = -i i^4 = 1 i^5 = i i^6 = -1 i^7 = -i See a pattern yet?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

1st step:  And now look at ASNinja's pattern